# COMMUNITYNOTES: A Dataset for Exploring the Helpfulness of Fact-Checking Explanations

### v1

@@ 4 Enhancing Helpfulness Prediction with Reason Definitions
Preliminary analysis shows that predicting why a particular explanation is (un)helpful is often subjective across instances, largely due to the lack of unified or official definitions. This subjectivity poses a barrier to model performance and limits the explainability of community-based fact-checking platforms. Prior work has demonstrated that providing clear label definitions can significantly improves the classification performance (Gao et al., 2023; Peskine et al., 2023; Khatuya et al., 2025). Given the large number of notes with annotated reason available, we aim to automatically generate and optimize reason label definitions based on data. To this end, we introduce a pipeline that leverages Automatic Prompt Optimization (APO) to generate and improve helpfulness reason definitions, which are then integrated with notes for reason prediction. Since the majority of notes and definitions are written in English, we focus on English data for the remainder of our experiments. Figure˜3 illustrates the overall pipeline.
4.1 Automatically Optimizing Reason Definition
We generate and optimize reason definitions following the first two steps in Figure˜3: (1) Seed reason definition generation and (2) Reason definition optimization. In the first step, we follow Peskine et al. (2023) and randomly sample 40 instances for each category. We then prompt the GPT-4o model to generate candidate definitions (see Table˜13 in Appendix for detailed prompts). [^4] Since a single sample may have multiple reason labels, we perform an additional check to ensure that each sample appears only within the same category to encourage diversity.
In the second step, we employ PromptAgent (Wang et al., 2024), a Monte Carlo Tree Search (MCTS)-based automatic prompt optimization framework, to refine the initial definitions. PromptAgent adopts an iterative refinement process that emulates how human experts strategically craft superior prompts, implemented through an MCTS-based planning mechanism. Each refined prompt corresponds to a new state in the search space, obtained by the refiner agent model through feedback collected from a feedback agent model. PromptAgent relies on the MCTS state–action value function to estimate the potential reward of each state. MCTS iteratively performs four key operations—selection, expansion, simulation, and backpropagation—until a predefined number of iterations is reached. The highest-reward trajectory is then selected to produce the final refined prompt. We initialize the process with the seed definitions generated in step (1) and run PromptAgent using a definition-refinement feedback loop.
@@ 4.1 Automatically Optimizing Reason Definition
To evaluate the impact of incorporating generated definitions, we perform zero-shot prompting with Llama3.1-8B-Instruct and Mistral-7B-Instruct for both helpfulness and reason prediction. Table˜6 reports results for the vanilla setting, as well as when incorporating seed and optimized definitions. We observe substantial F1 improvements in both helpfulness and reason prediction when adding seed definitions (on average around 15%), with further gains after definition optimization. For Llama3.1-8B-Instruct, both seed and optimized definitions yield large improvements. Although Mistral-7B-Instruct starts stronger than Llama3.1-8B-Instruct in the vanilla setting, it still benefits notably from incorporating definitions, particularly in reason-level F1. These results suggest that incorporating reason definition information substantially enhances model performance—not only in predicting whether a given explanation note is helpful, but also in identifying the underlying reason with respect to the claim (See seed definition in Table˜17 and optimized reason definition in Appendix LABEL:tab:optimized_def).
4.2 Training with Reason Definition Fusion
After obtaining the optimized helpfulness reason definitions, we proceed to the third step in Figure˜3, where we explore how to incorporate this information into the predictor’s training process. The core idea is that, for a given note associated with a claim, the predictor should attend more strongly to the relevant helpfulness reasons. Since SLM performs better in reason prediction, we use SLM as our predictors in the experiments.
We first obtain note embeddings $E_{note}$ from the original predictors and reason embeddings $E_{reason}$ from the optimized reason definitions using LLM2Vec (BehnamGhader et al., 2024), an unsupervised LLM-based embedding model. [^5] Specifically, we use LLM2Vec-Meta-Llama-31-8B-Instruct-mntp-supervised. We then apply a Multi-Head Attention (MHA) mechanism, using $E_{note}$ as the query (Q) and $E_{reason}$ as the key (K) and value (V), to fuse the reason information into the note representation. Finally, we adopt a multi-task learning objective that jointly optimizes helpfulness prediction and multi-label reason classification.
Table˜7 shows the impact of integrating the optimized reason definitions into the base models using the MHA module. Across all predictors, incorporating reason definitions via MHA consistently improves both helpfulness and reason F1 scores. For instance, RoBERTa-large-MHA achieves a Helpfulness F1 of 0.899 and a Reason F1 of 0.669, outperforming its vanilla counterpart (0.890 and 0.663, respectively). Similar trends hold for ModernBERT and DeBERTa, with DeBERTa-large-MHA achieving the highest Reason F1 score of 0.677. These results demonstrate that attention-based fusion of reason definitions effectively enhances SLMs, improving their ability to predict both whether a note is helpful and the underlying reason.
@@ 4.2 Training with Reason Definition Fusion
Table 7: Performance comparison of base models versus fusing optimized reason definition with the Multi-Head Attention module.
5 Generalization of the Explanation Helpfulness
In this section, we assess the generalization and the utility of our helpfulness predictors beyond community notes. A key challenge is that most existing datasets lack explicit annotations for explanation of helpfulness. Therefore, we explore two key scenarios in automated fact-checking. First, we evaluate whether helpfulness prediction can generalize to evidence sufficiency prediction, which is an important subtask in fact-checking. Second, we explore the effect of incorporating evidence helpfulness in automated fact-checking.
5.1 Generalization to Evidence Sufficiency
Evidence sufficiency is a task that decides whether the information in the evidence is sufficient for predicting the veracity of a claim. Automated fact-checking models can make veracity predictions only when there is enough evidence (Atanasova et al., 2022). In this scenario, we would like to explore whether our helpfulness predictors can predict the sufficiency of an evidence (since a helpful evidence must be “sufficient”). We adopted Sufficient Fact (Atanasova et al., 2022), a diagnostic fact-checking dataset to detect when evidence with omitted information is (in)sufficient. To adapt our prediction, we treat sufficient evidence (EI-ENOUGH INFO) as helpful and insufficient evidence (NEI-NOT ENOUGH INFO) as unhelpful.
@@ 5.1 Generalization to Evidence Sufficiency
Table˜8 shows the model performance on the evidence sufficiency task, focusing on the NEI label. As expected, fine-tuned (FT) models outperform the original notes predictors, with RoBERTa-large-FT achieving the best NEI F1 of 0.908. Nevertheless, the notes models show competitive performance: for example, ModernBERT-large-notes achieves a NEI F1 of 0.808, only about 10% lower than its FT counterpart. Across all predictors, the top predicted reason is notHelpfulMissingKeyPoints, which closely corresponds to the NEI label, as both signify the absence of critical information.
This indicates that helpfulness prediction carries transferable signal for evidence sufficiency. Still, helpfulness and sufficiency are not the same: even sufficient evidence could be predicted as unhelpful for other reasons such as notHelpfulArgumentativeOrBiased, which may account for the observed performance gap.
@@ 5.2 Incorporating Evidence Helpfulness in Automated Fact Checking
In this part, we further investigate whether predicted evidence helpfulness can enhance automated fact-checking. Helpfulness scores provide auxiliary signals for assessing evidence quality, with the potential to help models prioritize informative evidence during prediction. We adopted CLIMATE-FEVER (Diggelmann et al., 2020), a real-world dataset for verifying climate change–related claims. Each claim is associated with multiple evidence sentences, to which we apply our predictors to assign helpfulness scores and reasons. To perform fact-checking, we use Llama3.1-8B-Instruct to perform fact-checking (See detailed prompt in Appendix LABEL:tab:fc_prompt) prompting the model with both the claim and its associated evidence.
We compare the F1 with and without the helpfulness information of evidence and report results in Table˜9. We can see that, in general, incorporating helpfulness improves fact-checking performance with ModernBERT-base achieved an accuracy of 0.537. ModernBERT and RoBERTa-base improve fact-checking performance by around 2%. These findings indicate that providing auxiliary evidence helpfulness and reasons can benefit automated fact-checking. While ModernBERT demonstrates major gains, DeBERTa and RoBERTa yield smaller, non-significant improvements. Interestingly, this pattern mirrors their performance drop in cross-domain evaluation of evidence sufficiency in Section˜5.1 (despite having the best in-domain results in Table˜5). One possible explanation could be that these models have overfitted for the Community Notes domain and as such are less effective in other domains.
@@
Table 9: Automated fact-checking performance of Llama3.1-8B-instruct on the CLIMATE-FEVER dataset. F1 stands for Accuracy. +Helpfulness F1 means accuracy incorporating helpfulness information. Numbers in bold represents the gap between +Helpfulness and basic accuracy is statistically significant (p-value <0.05).
@@ 6 Conclusion
In this paper, we introduced the novel task of predicting the helpfulness of explanations and their underlying reasons for a given claim. To support this task, we constructed COMMUNITYNOTES, a large-scale post–note helpfulness dataset based on X’s Community Notes, and established benchmarks with pretrained language models. While existing models perform reasonably well on overall helpfulness, they struggle to identify underlying reasons. To address this, we proposed a framework that automatically generates and optimizes reason definitions, which, when reintegrated, greatly improves reason prediction. We further demonstrated that helpfulness prediction can help predict evidence sufficiency and enhance fact-checking systems.
We hope that our dataset and framework can serve as a foundation for building stronger helpfulness predictors capable of assigning prior labels to real-time community notes, thereby streamlining the community-based fact-checking process. Additionally, the automatically generated reason definitions may enhance the explainability and trustworthiness of fact-checking models and could be extended to improve other classification tasks beyond this domain.
@@
Our work aims to improve the transparency, explainability, and scalability of community-based fact-checking systems. By modeling and predicting the helpfulness of user-generated explanations, we hope to assist both contributors and platforms in identifying notes that are more informative, balanced, and useful for the public discourse. However, several ethical considerations deserve discussion.
@@
Community-based ratings may reflect social, cultural, or political biases that influence which notes are deemed “helpful.” Our models inevitably inherit such biases from the underlying data. We emphasize that model predictions should not be used to automatically moderate or suppress user content. Instead, they should be interpreted as supporting signals to guide human review.
@@
While our framework can aid in prioritizing or surfacing potentially helpful notes, it is not intended to replace human judgment. Automated systems that assign helpfulness or credibility labels at scale could have downstream impacts on content visibility, user reputation, and the dynamics of online discourse. We encourage future work to incorporate fairness-aware and interpretability-centered mechanisms before deployment.
@@
Despite these caveats, we believe our dataset and models can contribute positively to combating misinformation by improving the quality, trustworthiness, and accessibility of community fact-checking. By making the definitions of helpfulness explicit and machine-interpretable, we hope to support more transparent, accountable, and participatory approaches to information verification.
@@
All data used in this work originate from the publicly available X Community Notes dataset, [^6] https://communitynotes.x.com/guide/en/under-the-hood/download-data which is released by the platform under its open data policy. The dataset consists of user-generated notes, their associated posts, and corresponding helpfulness and reason labels. No personally identifiable information (PII), user handles, or private metadata are included in our version of the dataset. We processed and released only anonymized text and aggregated statistics in compliance with the platform’s terms of service. The dataset is multilingual, with English, Japanese, and Spanish as the most frequent languages. All examples are drawn from publicly visible content. Any future release of derived data or annotations will follow ACL’s data ethics guidelines and include proper documentation describing preprocessing steps, filtering criteria, and licensing terms.
@@
Our study does not introduce direct security risks, but the use of models trained on user-generated content entails potential vulnerabilities. Automated systems that evaluate or rank explanatory notes could be misused for large-scale manipulation—for example, to promote particular narratives or suppress dissenting perspectives. To mitigate such risks, we release only aggregated and anonymized data, and we emphasize that our models are intended for research and analysis, not for automated moderation or deployment without human oversight. Future work should include safeguards against adversarial attacks, bias exploitation, and model inversion risks, particularly when integrating such predictors into live social platforms.
@@ Appendix A Appendix
A.1 Extended Background of X Community Notes
@@
Community Notes originated from the Birdwatch (Wojcik et al., 2022) program launched by X (formerly Twitter) in 2021.777Introducing Birdwatch, a community-based approach to misinformation, By Keith Coleman, Monday, 25 January 2021 Participants, known as contributors, write “notes” that clarify or contextualize claims made in posts, and other users rate these notes for helpfulness and accuracy. Notes receiving broad agreement across diverse raters are surfaced publicly beneath the corresponding posts. This community-driven system aims to promote transparency and provide multiple perspectives in addressing misinformation on social media (Augenstein et al., 2025). Initially limited to U.S. users, the program gained wider attention in 2022 during major misinformation events such as the Russian invasion of Ukraine and the COVID-19 pandemic. Birdwatch was re-branded as Community Notes and expanded globally in November 2022. As of November 2023, the program had over 133,000 contributors, with notes reportedly viewed tens of millions of times per day, reflecting its growing role in mitigating misinformation online.
@@
Community Notes introduces a promising new paradigm for fact-checking. It democratizes content moderation by enabling the public to collectively determine whether a note should be added to a post, offering a less intrusive alternative to traditional expert-based fact-checking systems that often rely on explicit warning labels. This community-driven approach also accelerates the detection of misinformation, offering better scalability. However, Community Notes also faces several challenges. The community can still favor popularity over truth, similar to the echo-chamber effects often seen on social media. While its scalability and faster operation are promising, studies show that its performance still lags behind expert-based fact-checking in many cases (Martel and Rand, 2024). In addition, helpful notes often appear too late—after a post has already received significant engagement—reducing their effectiveness in correcting misinformation and changing user beliefs (Augenstein et al., 2025; De et al., 2025).

### v2

@@ 4 Enhancing Helpfulness Prediction with Reason Definitions
Preliminary analysis shows that predicting why a particular explanation is (un)helpful is often subjective across instances, largely due to the lack of unified or official definitions. This subjectivity poses a barrier to model performance and limits the explainability of community-based fact-checking platforms. Prior work has demonstrated that providing clear label definitions can significantly improves the classification performance (Gao et al., 2023; Peskine et al., 2023; Khatuya et al., 2025).
Given the large number of notes with annotated reasons available, we aim to automatically generate and optimize reason label definitions based on data. To this end, we introduce a pipeline that leverages Automatic Prompt Optimization (APO) to generate and improve helpfulness reason definitions, which are then integrated with notes for reason prediction. Since the majority of notes are written in English, we focus on English data for the remainder of our experiments. Figure˜5 illustrates the overall pipeline of our approach.
4.1 Automatically Optimizing Helpfulness Reason Definition
We automatically generate and optimize reason definitions following the first two steps in Figure˜5: (1) Seed reason definition generation; and (2) Reason definition optimization. In the first step, we follow Peskine et al. (2023) and randomly sample 40 instances per helpfulness reason category. We then prompt the GPT-4o model [^4] Detailed model version: gpt-4o-2024-08-06. to generate an initial “seed” version of the definitions based on the sampled candidate instances (see Appendix LABEL:tab:gen_def for detailed prompts to generate seed definitions). [^5] Since a single sample may have multiple reason labels, we perform an additional check to ensure that each sample appears only within the same category to encourage diversity. In the second step, we employ PromptAgent (Wang et al., 2024), a Monte Carlo Tree Search (MCTS)-based automatic prompt optimization framework, to refine the initial definitions.
PromptAgent adopts an iterative refinement process that emulates how human experts strategically craft effective prompts, using an MCTS-based planning mechanism. Each refined prompt corresponds to a new state in the search space, obtained by the refiner agent model through feedback collected from the feedback agent model. PromptAgent relies on the MCTS state–action value function to estimate the reward of each state. MCTS iteratively performs four key operations—selection, expansion, simulation, and backpropagation—until a predefined number of iterations is reached. The highest-reward trajectory is then selected to produce the final refined prompt. We initialize the process with the seed definitions obtained from step (1) in Figure˜5 and run adapted PromptAgent using a definition-refinement feedback loop.
@@ 4.1 Automatically Optimizing Helpfulness Reason Definition
To evaluate the impact of incorporating generated definitions, we perform zero-shot prompting with Llama3.1-8B and Mistral-7B for both helpfulness and reason prediction. Table˜6 reports results for the vanilla setting, as well as for models integrated with seed and optimized definitions. We observe substantial F1 improvements in both helpfulness and reason prediction when adding seed definitions (on average around 15%), with further gains after definition optimization, indicating that refined definitions provide additional benefits.
For Llama3.1-8B, both seed and optimized definitions yield large improvements. Although Mistral-7B starts stronger than Llama3.1-8B in the vanilla setting, it still benefits notably from incorporating definitions, particularly in reason F1. These results suggest that incorporating reason definition information substantially enhances model performance—not only in predicting whether a given explanation note is helpful, but also in identifying the underlying reason with respect to the claim. (See detailed seed definition in LABEL:tab:seed_def and optimized reason definition in LABEL:tab:optimized_def in the Appendix).
@@
Table 7: Reason prediction performance comparison of base SLMs versus fusing optimized reason definition with the Multi-Head Attention module.
4.2 Training with Reason Definition Fusion
After obtaining the optimized helpfulness reason definitions, we proceed to the third step in Figure˜5, where we incorporate this information into the training SLM predictors for helpfulness and reason prediction. The core idea is that the predictor should attend more strongly to the most relevant helpfulness reasons for a given claim–note pair.
We obtain note embeddings $E_{note}$ from the base predictors and reason embeddings $E_{reason}$ from the optimized reason definitions using LLM2Vec (BehnamGhader et al., 2024). [^6] Specifically, we use the following model: LLM2Vec-Meta-Llama-31-8B-Instruct-mntp-supervised. We then apply a Multi-Head Attention (MHA) mechanism, using $E_{note}$ as the query (Q) and $E_{reason}$ as the key (K) and value (V), to fuse the reason information into the note representation. Finally, we adopt a multi-task learning objective that jointly optimizes helpfulness prediction and multi-label reason classification for our experiments.
Table˜7 shows that integrating optimized reason definitions via MHA consistently improves performance across all SLMs. For example, RoBERTa-large-MHA achieves a Helpfulness F1 of 0.899 and a Reason F1 of 0.669, outperforming its vanilla counterpart (0.890 and 0.663, respectively). Similar trends hold for ModernBERT and DeBERTa, with DeBERTa-large-MHA achieving the highest Reason F1 score of 0.677.
Overall, these results demonstrate that attention-based fusion of reason definitions effectively enhances SLMs for both helpfulness and reason prediction on COMMUNITYNOTES.
@@ 4.2 Training with Reason Definition Fusion
5 Generalization of the Explanation Helpfulness
In this section, we assess the generalization and the utility of our helpfulness predictors beyond community notes. A key challenge is that most existing datasets lack explicit annotations for explanation of helpfulness. Therefore, we explore two key scenarios in automated fact-checking. First, we evaluate whether helpfulness prediction can generalize to evidence sufficiency prediction, which is an important subtask in fact-checking (Atanasova et al., 2022). Second, we explore the effect of incorporating evidence helpfulness in automated fact-checking (Xing et al., 2025).
5.1 Generalization to Evidence Sufficiency
Evidence sufficiency is a task that decides whether the information in the evidence is sufficient for predicting the veracity of a claim. Automated fact-checking models can make veracity predictions only when there is enough evidence (Atanasova et al., 2022). In this scenario, we would like to explore whether our helpfulness predictors can predict the sufficiency of an evidence (since a helpful evidence must be “sufficient”). We adopted Sufficient Fact (Atanasova et al., 2022), a diagnostic fact-checking dataset to detect when evidence with omitted information is (in)sufficient. To adapt our prediction, we treat sufficient evidence (EI-ENOUGH INFO) as helpful and insufficient evidence (NEI-NOT ENOUGH INFO) as unhelpful. Table˜8 shows the model performance on the evidence sufficiency task, focusing on the NEI label.
As expected, fine-tuned (FT) models outperform the original notes predictors, with RoBERTa-large-FT achieving the best NEI F1 of 0.908. Nevertheless, the notes models show competitive performance: for example, ModernBERT-large-notes achieves a NEI F1 of 0.808, only about 10% lower than its FT counterpart. Across all predictors, the top predicted reason is notHelpfulMissingKeyPoints, which closely corresponds to the NEI label, as both signify the absence of critical information. This indicates that helpfulness prediction carries transferable signal for evidence sufficiency.
Still, helpfulness and sufficiency are not the same: even sufficient evidence could be predicted as unhelpful for other reasons such as notHelpfulArgumentativeOrBiased, which may account for the observed performance gap in our experiments.
@@ 5.2 Incorporating Evidence Helpfulness in Automated Fact Checking
In this part, we further investigate whether predicted evidence helpfulness can enhance automated fact-checking. Helpfulness scores provide auxiliary signals for assessing evidence quality, with the potential to help models prioritize informative evidence during prediction. We adopted CLIMATE-FEVER (Diggelmann et al., 2020), a real-world dataset for verifying climate change–related claims. Each claim is associated with multiple evidence sentences, to which we apply our predictors to assign helpfulness scores and reasons. To perform fact-checking, we use Llama3.1-8B to perform fact-checking (see detailed prompt in Appendix LABEL:tab:fc_prompt) prompting the model with both the claim and its associated evidence.
We compare the F1 with and without the evidence helpfulness information and report results in Table˜9. Overall, incorporating helpfulness improves fact-checking performance, with ModernBERT-base achieving 0.537 accuracy and ModernBERT and RoBERTa-base improving performance by around 2%. While ModernBERT demonstrates major gains, DeBERTa and RoBERTa yield smaller, non-significant improvements. Interestingly, this pattern mirrors their performance drop in cross-domain evidence sufficiency evaluation in Section˜5.1 (despite strong in-domain results in Table˜5). One possible explanation is that models overfit COMMUNITYNOTES and generalize less effectively to other domains.
@@
Table 9: Automated fact-checking performance of Llama3.1-8B on the CLIMATE-FEVER dataset. F1 stands for Accuracy. +Helpfulness F1 means accuracy incorporating helpfulness information. Numbers in bold represents the gap between +Helpfulness and basic prediction accuracy is statistically significant (using a two-tailed t-test with $p<0.05$).
@@ 6 Conclusion
In this paper, we introduced the novel task of predicting the helpfulness of explanations and their underlying reasons for a given claim. To support this task, we constructed COMMUNITYNOTES, a large-scale post–note helpfulness dataset based on X’s Community Notes, and established benchmarks with pretrained language models.
While existing models perform reasonably well on overall helpfulness, they struggle to identify underlying reasons. To address this limitation, we proposed a framework that automatically generates and optimizes reason definitions, which, when reintegrated, greatly improves reason prediction. We further demonstrated that helpfulness prediction can help predict evidence sufficiency and enhance automated fact-checking systems.
Beyond performance gains, our findings highlight the importance of making evaluation criteria for explanation helpfulness more explicit and structured. We hope that our dataset and framework can serve as a foundation for building stronger helpfulness predictors capable of assigning prior labels to real-time community notes, thereby streamlining the community-based fact-checking process. Additionally, the automatically generated and optimized reason definitions may enhance the explainability and trustworthiness of automated fact-checking models and could be extended to improve other classification tasks beyond this domain.
In future work, we plan to extend our framework to multilingual and multimodal settings and evaluate it in more advanced end-to-end fact-checking systems and real-world applications.
@@
Our work aims to improve the transparency, explainability, and scalability of community-based fact-checking systems. By modeling and predicting the helpfulness of user-generated explanations, we hope to assist both contributors and platforms in identifying notes that are more informative, balanced, and useful for the public discourse.
@@
Community-based ratings may reflect social, cultural, or political biases that influence which notes are deemed “helpful.” Our models inevitably inherit such biases from the underlying data. We would like to emphasize that model predictions should not be directly used to automatically moderate or suppress user content. Instead, they should be interpreted as auxiliary signals to guide human review.
@@
While our framework can help prioritize potentially helpful notes, it is not intended to replace human judgment. Automated systems that assign helpfulness or credibility labels at scale may affect content visibility, user reputation, and online discourse. We encourage incorporating fairness-aware and interpretable mechanisms before deployment.
@@
Despite these caveats, our dataset and models can help combat misinformation by improving the quality, trustworthiness, and accessibility of community fact-checking. Making helpfulness definitions explicit and machine-interpretable supports more transparent and accountable information verification.
@@
All data that we used in this work originated from the publicly available X Community Notes dataset, [^7] https://communitynotes.x.com/guide/en/under-the-hood/download-data which was released under the X platform’s open data policy. The dataset contains user-generated notes, associated posts, and helpfulness and reason labels, without any personally identifiable information. We release only anonymized text and aggregated statistics in compliance with the X platform’s terms of service. The dataset is multilingual, with English, Japanese, and Spanish being the best represented languages. Any future release of derived data will follow ACL data ethics guidelines with appropriate documentation, to enable transparent, reproducible research in NLP and computational social science.
@@
Our study introduces no direct security risks, but models trained on user-generated content may be vulnerable to misuse, such as large-scale manipulation or suppression of dissent. To mitigate these risks, we release only aggregated and anonymized data and emphasize that our models are intended for research use, not automated moderation or deployment without human oversight. Future work should address adversarial attacks, bias exploitation, and model inversion risks when integrating such models into live platforms.
Acknowledgments
We thank our reviewers for their valuable reviews and feedback, which greatly contributed to the improvement of our paper.
@@ Appendix A Appendix
A.1 Extended Background of X Community Notes Program
@@
Community Notes originated from the Birdwatch (Wojcik et al., 2022) program which was launched by X (formerly Twitter) in 2021.888Introducing Birdwatch, a community-based approach to misinformation, By Keith Coleman, Monday, 25 January 2021. Participants, known as contributors, write “notes” that clarify or contextualize claims made in posts, and other users rate these notes for helpfulness and accuracy. Notes receiving broad agreement across diverse raters are surfaced publicly beneath the corresponding posts. This community-driven system aims to promote transparency and provide multiple perspectives in addressing misinformation on social media (Augenstein et al., 2025). Initially limited to U.S. users, the program gained wider attention in 2022 during major misinformation events such as the Russian invasion of Ukraine and the COVID-19 pandemic. Birdwatch was re-branded as Community Notes and expanded globally in November 2022. As of November 2023, the program had over 133,000 contributors, with notes reportedly viewed tens of millions of times per day, reflecting its growing role in mitigating misinformation in online social media.
@@
Community Notes introduces a promising new paradigm for fact-checking. It democratizes content moderation by enabling the public to collectively determine whether a note should be added to a post, offering a less intrusive alternative to traditional expert-based fact-checking systems that often rely on explicit warning labels. This community-driven approach also accelerates the detection of misinformation, offering better scalability. However, Community Notes also faces several challenges. The community can still favor popularity over truth, similar to the echo-chamber effects often seen on social media. While its scalability and faster operation are promising, studies show that its performance still lags behind expert-based fact-checking in many cases (Martel and Rand, 2024). In addition, helpful notes often appear too late—after a post has already received significant engagement—reducing their effectiveness in correcting misinformation and changing user beliefs (Augenstein et al., 2025; De et al., 2025; Borenstein et al., 2025).
Phase # Step Description
Prescoring 1. Pre-filter Data: Include only raters ($\geq 10$ ratings) & notes ($\geq 5$ ratings). Coalesce similar raters.
4. Refine Score: Fit a tag-consensus MF model on the helpfulness-filtered data and update all helpfulness scores.
3. Fit Diligence Model: Fit the note diligence MF model.
6. Assign Explanation Tags: Assign the top two matching explanation tags. If two such tags are not found, revert status to NEED_MORE_RATINGS.
Table 10: Overview of the X Community Notes note ranking Algorithm pipeline.